Let f(x)=4-x^ and g(x)=2-x. Find f/g(x).
Answer by stanbon(75887)

Let f(x)=4-x^2 and g(x)=2-x. Find f/g(x).
---------
(f/g)x = (f(x))/(g(x))
= (4-x^2)/(2-x)
= [(2-x)(2+x)/(2-x)]
= 2+x
==================
Cheers,
Stan H.





==========
Cheers,
Stan H.